OpenStudy (anonymous):

simplify the expression and assume that all variables are positive (k^1/2)^-7/(k^10)^1/4

OpenStudy (anonymous):

Do you know how to simplify the numerator and denominator separately? (power-to-a-power, multiply exponents)

OpenStudy (anonymous):

the closest answer I have is 1/k^7 not sure what I am getting wrong

OpenStudy (anonymous):

That's close; might be a subtraction error.

OpenStudy (anonymous):

Are you getting \[\frac{k^{-7/2}}{k^{5/2}}?\]

OpenStudy (zzr0ck3r):

(k^(-7/2))/k^(5/2) = k^(-7/2-5/2) = k^(-6)

OpenStudy (zzr0ck3r):

= 1/k^6